# Theoretical framework for lattice QCD computations of $B\to K\ell^{+}\ell^{-}$ and $\bar{B}_{s}\to\ell^{+}\ell^{-}\gamma$ decays rates, including contributions from “Charming Penguins”

### v1

@@ Abstract
We develop a strategy for computing the $B\to K\ell^{+}\ell^{-}$ and $\bar{B}_{s}\to\gamma\ell^{+}\ell^{-}$ decay amplitudes using lattice QCD (where $\ell^{\pm}$ are charged leptons). We focus on those terms which contain complex contributions to the amplitude, due to on-shell intermediate states propagating between the weak operator and electromagnetic current(s). Such terms, which are generally estimated using model calculations and represent significant uncertainties in the phenomenological predictions for these decays, cannot be computed using standard lattice QCD techniques. It has recently been shown that such contributions can be computed using spectral-density methods and our proposed strategy, which we discuss in detail, is built on this approach. The complex contributions include the “charming penguins” (matrix elements of the current-current operators $O_{1}^{(c)}$ and $O_{2}^{(c)}$ defined in Eq. (6) below), in which the charm-quark loop can propagate long distances, particularly close to the region of charmonium resonances. They also include the contributions from the chromomagnetic operator ($O_{8}$ in standard notation, defined in Eq. (8) below). We discuss the renormalization of the ultra-violet divergences, and in particular those which arise due to “contact” terms, and explain how those which appear as inverse powers of the lattice spacing can be subtracted non-perturbatively. We apply the spectral density methods in an instructive exploratory computation of the charming penguin diagram in $B\to K\ell^{+}\ell^{-}$ decays in which the virtual photon is emitted from the charm-quark loop (the diagram in Fig. 1(a) below) and discuss the prospects and strategies for the reliable determination of the amplitudes in future dedicated computations.
@@ I Introduction
In the Standard Model, Flavor Changing Neutral Current (FCNC) $B_{(s)}$ decays are strongly suppressed and thus represent a window for discovering New Physics (NP) effects. At the quark level $b\to s(d)$ transitions give rise to a variety of interesting processes that have been measured experimentally and compared to theoretical predictions, e.g $B\to K^{*}\gamma$ or $B\to K^{(*)}\ell^{+}\ell^{-}$ and $\bar{B}_{s}\to\gamma\ell^{+}\ell^{-}$[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] . The main limitation in the comparison between experimental measurements and theoretical predictions is that the evaluation of the hadronic matrix elements generally relies on phenomenological models for which it is very difficult to estimate the uncertainty. In this paper we show that, as the result of a number of recent theoretical developments [23, 24], a full class of semileptonic FCNC decays, including several of the ones mentioned above, can be computed quantitatively from first principles using Lattice QCD, paving the way towards truly quantitative, and in the future precise, tests of the theory. The precise measurements of $B\to K^{(*)}\ell^{+}\ell^{-}$ decays by LHCb [10, 11, 13, 25, 26, 27], see also [7, 8, 9, 12, 14, 15, 16, 17, 19], opened a new exciting sector of FCNC phenomenology. While the hints of the presence of NP in the ratio of branching fractions $R_{K}=B(B^{+}\to K^{+}\mu^{+}\mu^{-})/B(B^{+}\to K^{+}e^{+}e^{-})$ [28, 29], and similarly in $R_{K^{*}}$, $R_{K_{S}}$, $R_{K^{*+}}$ [30, 31, 32], have disappeared [33, 34], there nevertheless remain some tensions between SM estimates and experimental measurements in absolute $BR$s and angular distributions [35, 36, 37, 38, 39, 40, 41] (see however [42, 43, 44, 45]), and it is therefore very important to produce reliable theoretical predictions of these quantities in the SM and beyond. In this paper we show how this can be achieved for $B\to K\ell^{+}\ell^{-}$ decays, and in particular how the charming penguin contributions, i.e. contributions from diagrams containing charm-quark loops, can be determined (see, for example, Fig. 1).
Figure 1: Connected quark-flow diagrams for the process $B\to K\ell^{+}\ell^{-}$ which contain a charm-quark loop. The shaded circle, marked $O_{1,2}^{(c)}$, represents either of the two current-current operators $O_{1}^{(c)}$ or $O_{2}^{(c)}$ defined in Eq. (6). The charged leptons $\ell^{\pm}$, which couple to the virtual photon, $\gamma^{\ast}$, are not shown.
@@
The decay $B\to K^{\ast}\gamma$ is, as yet, an unexplored channel in lattice calculations and presents additional issues related to the instability of the $K^{\ast}$ meson. In particular the reconstruction of the Minkowski amplitude from Euclidean lattice correlation functions is more complex and the necessary theoretical framework has not yet been fully developed. For this reason we do not discuss $B\to K^{\ast}\ell^{+}\ell^{-}$ and $B\to K^{\ast}\gamma$ decays further in this paper.
For real physical quantities, such as leptonic decay constants, semileptonic form factors (between states which are stable under strong interactions) and the $B$-parameters of neutral pseudoscalar-meson mixing, the continuation from Minkowski to Euclidean space poses no difficulty and they can readily be determined from Lattice QCD computations of Euclidean correlation functions (see, for example, the comprehensive compilation of results in Ref. [47]). Since Euclidean correlation functions are real, difficulties in the continuation from Minkowski to Euclidean space arise when the physical quantities of interest are complex. This is particularly severe for the evaluation of the amplitudes for $B$-decays, where a large number of on-shell intermediate states can propagate and contribute to the imaginary part. In an attempt to overcome this complication, methods based on the spectral representation have been developed [48, 49, 50, 51, 23, 52, 53, 24, 54] and applied to a variety of physical processes, including Refs. [55, 56, 57, 58, 59, 60, 61, 62, 46, 63, 64, 65, 66, 67]. In particular, in Ref. [24] a method has been proposed which allows for the determination of both the real and imaginary parts of complex amplitudes in which the hadronic factor is of the generic form:
@@ III.2 Time-ordered diagrams contributing to $\bar{B}_{s}\to\gamma\ell^{+}\ell^{-}$ decay amplitudes
Figure 2: Schematic illustrations of the two time-orderings in Eq. (11) contributing to $B\to K\ell^{+}\ell^{-}$ decays. The dashed lines between the operators represent all propagating states with the indicated $B$ and $S$ quantum numbers. The emission of the virtual photon is not shown.
@@
Figure 3: Schematic illustrations of the three time-orderings in Eq. (14) which contribute an imaginary part to the $B\to\gamma\ell^{+}\ell^{-}$ decay amplitude. The dashed lines between the operators represent all propagating states with the indicated $B$ quantum number. The emission of the real and virtual photons is not shown.
@@ III.3 Generalized spectral density representation
where the subscript $E$ on the left-hand side denotes Euclidean. The HLT method [23] is based on the expansion of the smearing kernel in terms of exponentials as follows:
@@ VII.1 The contact terms
The term $H^{\nu+;\,3\mathrm{subs}}_{1,2}(\vec{q},\epsilon)$, defined in Eq. (77), has been designed so that it has no ultraviolet divergences, even at finite-values of $\epsilon$. We envisage applying the HLT procedure to each of the four terms in Eq. (77). The first three terms on the right-hand side of Eq. (76) compensate for the subtractions introduced in the definition of $H^{\nu+;\,3\mathrm{subs}}_{1,2}(\vec{q},\epsilon)$ given in Eq. (77). They have no poles in the range of the integration over $E$ so that we can set $\epsilon=0$ in these terms. They do however, have quadratic and linear ultraviolet divergences. In order to cancel the power divergences, we combine these three terms with the contribution from the region $t<0$ in Eq. (IV) in a correlated way so that only the logarithmic divergence remains. In a lattice QCD computation this combination is performed under the jackknife, with the expectation that the correlations between the four terms will reduce the resulting uncertainty.
@@
Figure 8: Diagrams for the decay $\bar{B}_{s}\to\mu^{+}\mu^{-}\gamma$ in which both the real photon $\gamma$ with momentum $k$ and the virtual photon with momentum $q$ are emitted from the charm-quark loop.
A corresponding procedure needs to be developed and applied to the contact terms present in the amplitude for $\bar{B}_{s}\to\gamma\ell^{+}\ell^{-}$ decays. We postpone the presentation of the details to a future publication, when we will be in a position to have learned from the experience of computing the charming penguin contributions to $B\to K\ell^{+}\ell^{-}$ decay amplitudes. Here we only note that for the process $\bar{B}_{s}\to\gamma\ell^{+}\ell^{-}$, there are also diagrams in which both the real and virtual photons are emitted from the charm-quark loop as in Fig. 8. Although the superficial degree of divergence of such diagrams would allow for a logarithmic ultraviolet divergence, and indeed, even at lowest order, each of the two diagrams in Fig. 8 separately has such a divergence, electromagnetic gauge invariance ensures that sum of the two contributions from these diagrams is finite. The requirement that contracting the matrix element in Eq. 50 with $k^{\mu}$ and $q^{\nu}=(p_{\bar{B}_{s}}-k)^{\nu}$ vanishes, implies that the matrix element is proportional to two powers of the external momenta which is sufficient to make it finite. However, similarly to the above discussion of the cancellation of the power divergences in $B\to K\ell^{+}\ell^{-}$ decays, the cancellation of the logarithmic ultraviolet divergences in the diagrams of Fig.8 requires the sum of the contributions from all the time-orderings. This is true since we use the lattice conserved vector current. Had we used the local vector current a more complicated subtraction procedure would need to be implemented but we do not discuss this here.
@@ VII.2.4 Renormalization of the operator $O_{8}$
Figure 9: Top Panel: Bare correlators $\widetilde{C}^{3}_{1,2;VV,AA}(t,\vec{q}\hskip 1.2pt)$, defined in Eqs. (129)-(130), corresponding to the $B\to K\ell^{+}\ell^{-}$ decay. Bottom: The correlators $\widetilde{C}^{3}_{1,2;VV,AA}(t,\vec{q}\hskip 1.2pt)$ corresponding to the $B\to\eta_{ss^{\prime}}\ell^{+}\ell^{-}$ decay, obtained by replacing the (spectator) light-quark in the $B\to K\ell^{+}\ell^{-}$ correlation function by a strange quark as described in the text. $N=1$ and $N=2$ correspond to the operators $O_{1}^{(c)}$ and $O_{2}^{(c)}$ respectively.
@@
$\widetilde{C}_{1,2;VV,AA}^{\nu}(t,\vec{q}\hskip 1.2pt)=\int_{E^{*}}^{\infty}\frac{dE}{2\pi}e^{-Et}\rho^{\nu+}_{1,2;VV,AA}(E,\vec{q}\hskip 1.2pt)~,\qquad\rho^{\nu+}_{1,2}(t,\vec{q}\hskip 1.2pt)=\rho^{\nu 3+}_{1,2;VV}(t,\vec{q}\hskip 1.2pt)+\rho^{\nu+}_{1,2;AA}(t,\vec{q}\hskip 1.2pt)~.$ (131)
@@ VIII.1 From the Euclidean correlators to the amplitude using the SFR/HLT method
where $\rm Cov$} is the covariance matrix of the Euclidean lattice correlator $\widetilde{C}_{1,2}^{\nu+}(t,\vec{q}\hskip 1.2pt;m_{H})$, and $\lambda$ is the so-called trade-off parameter. In Eq. (139), $\beta$ and $E_{\mathrm{th}}$ are algorithmic parameters which we set to $\beta=2^{-}$ and $E_{\rm th}=3.5\,{\rm GeV}$ ($3.7~{\rm GeV}$) in the case of $B\to K\ell^{+}\ell^{-}$ ($B_{s}\to\eta_{ss^{\prime}}\ell^{+}\ell^{-}$). [^13] For the convergence of the integral $\beta$ must be smaller than 2. In practice we set its value to $1.99$. In the presence of statistical errors, the functional $B$ disfavours coefficients $\bm{g}$ leading to large statistical uncertainties in the reconstructed smeared amplitude. The balance between having small systematic errors (small $A_{R}^{\beta}[\bm{g}]$ and $A_{I}^{\beta}[\bm{g}]$) and small statistical errors (small $B[\bm{g}]$) depends on the tunable parameter $\lambda$. Its optimal value $\lambda^{\rm opt}$ is determined by monitoring the evolution of the reconstructed smeared amplitude for different values of $\lambda$. The optimal value is chosen to be in the statistically-dominated regime, where $\lambda$ is sufficiently small that the systematic error due to the kernel reconstruction is smaller than the statistical error (therefore, in this region, the results are stable under variations of $\lambda$ within statistical uncertainties), but still large enough to have reasonably small statistical uncertainties. The kernel functions that we use at non-zero lattice spacing are obtained from Eq. (VIII.1) by the replacement

### v2

# Theoretical framework for lattice QCD computations of $B\to K\ell^{+}\ell^{-}$ and $\bar{B}_{s}\to\ell^{+}\ell^{-}\gamma$ decays rates, including contributions from charming penguin diagrams

@@ Abstract
We develop a strategy for computing the $B\to K\ell^{+}\ell^{-}$ and $\bar{B}_{s}\to\gamma\ell^{+}\ell^{-}$ decay amplitudes using lattice QCD (where $\ell^{\pm}$ are charged leptons). We focus on those terms which contain complex contributions to the amplitude, due to on-shell intermediate states propagating between the weak operator and electromagnetic current(s). Such terms, which are generally estimated using model calculations and represent significant uncertainties in the phenomenological predictions for these decays, cannot be computed using standard lattice QCD techniques. It has recently been shown that such contributions can be computed using spectral-density methods and our proposed strategy, which we discuss in detail, is built on this approach. The complex contributions include the “charming penguins” (matrix elements of the current-current operators $O_{1}^{(c)}$ and $O_{2}^{(c)}$ defined in Eq. (6) below), in which the charm-quark loop can propagate long distances, particularly close to the region of charmonium resonances. They also include the contributions from the chromomagnetic operator ($O_{8}$ in standard notation, defined in Eq. (8) below). We discuss the renormalization of the ultra-violet divergences, and in particular those which arise due to “contact” terms, and explain how those which appear as inverse powers of the lattice spacing can be subtracted non-perturbatively. We apply the spectral density methods in an instructive exploratory computation of the charming penguin diagram in $B\to K\ell^{+}\ell^{-}$ decays in which the virtual photon is emitted from the charm-quark loop (the diagram in Fig. 1(a) below) and discuss the prospects and strategies for the reliable determination of the amplitudes in future dedicated computations; computations which are however, beyond the scope of the present paper.
@@ I Introduction
In the Standard Model, Flavor Changing Neutral Current (FCNC) $B_{(s)}$ decays are strongly suppressed and thus represent a window for discovering New Physics (NP) effects. At the quark level $b\to s(d)$ transitions give rise to a variety of interesting processes that have been measured experimentally and compared to theoretical predictions, e.g $B\to K^{*}\gamma$ or $B\to K^{(*)}\ell^{+}\ell^{-}$ and $\bar{B}_{s}\to\gamma\ell^{+}\ell^{-}$[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] . The main limitation in the comparison between experimental measurements and theoretical predictions is that the evaluation of the hadronic matrix elements generally relies on phenomenological models for which it is very difficult to estimate the uncertainty. In this paper we show that, as the result of a number of recent theoretical developments [23, 24], a full class of semileptonic FCNC decays, specifically those with 0 or 1 stable hadron in the final state and hence including several of the ones mentioned above, can be computed quantitatively from first principles using Lattice QCD, paving the way towards truly quantitative, and in the future precise, tests of the theory.
We start by explaining the phenomenological motivation for developing the framework by briefly reviewing the experimental status of $b\to s$ exclusive measurements. The precise measurements of $B\to K^{(*)}\ell^{+}\ell^{-}$ decays by LHCb [10, 11, 13, 25, 26, 27], see also [7, 8, 9, 12, 14, 15, 16, 17, 19], opened a new exciting sector of FCNC phenomenology. While the hints of the presence of NP in the ratio of branching fractions $R_{K}=B(B^{+}\to K^{+}\mu^{+}\mu^{-})/B(B^{+}\to K^{+}e^{+}e^{-})$ [28, 29], and similarly in $R_{K^{*}}$, $R_{K_{S}}$, $R_{K^{*+}}$ [30, 31, 32], have disappeared [33, 34], there nevertheless remain some tensions between SM estimates and experimental measurements in absolute $BR$s and angular distributions [35, 36, 37, 38, 39, 40, 41, 42] (see however [43, 44, 45, 46]), and it is therefore very important to produce reliable theoretical predictions of these quantities in the SM and beyond. In this paper we show how this can be achieved for $B\to K\ell^{+}\ell^{-}$ decays, and in particular how the charming penguin contributions, i.e. contributions from diagrams containing charm-quark loops, can be determined (see, for example, Fig. 1).
Figure 1: Connected quark-flow diagrams for the process $B\to K\ell^{+}\ell^{-}$ which contain a charm-quark loop. The shaded circle, marked $O_{1,2}^{(c)}$, represents either of the two current-current operators $O_{1}^{(c)}$ or $O_{2}^{(c)}$ defined in Eq. (6). The charged leptons $\ell^{\pm}$, which couple to the virtual photon, $\gamma^{\ast}$, are not shown. Panels (a)-(d) correspond, respectively, to photon emission from the charm, strange, down, and bottom quarks.
@@
The decay $B\to K^{\ast}\gamma$ is, as yet, an unexplored channel in lattice calculations and presents additional issues related to the instability of the $K^{\ast}$ meson. In particular the reconstruction of the Minkowski amplitude from Euclidean lattice correlation functions is more complex and the necessary theoretical framework has not yet been fully developed. For this reason we do not discuss $B\to K^{\ast}\ell^{+}\ell^{-}$ and $B\to K^{\ast}\gamma$ decays further in this paper, although we anticipate that the method developed in this paper may eventually be an important element in the determination of the corresponding amplitudes.
Motivated by the challenge to obtain reliable theoretical results with which the experimental results outlined above can be compared, we now begin a discussion of the theory. For real physical quantities, such as leptonic decay constants, semileptonic form factors (between states which are stable under strong interactions) and the $B$-parameters of neutral pseudoscalar-meson mixing, the continuation from Minkowski to Euclidean space poses no difficulty and they can readily be determined from Lattice QCD computations of Euclidean correlation functions (see, for example, the comprehensive compilation of results in Ref. [48]). Since Euclidean correlation functions are real, difficulties in the continuation from Minkowski to Euclidean space arise when the physical quantities of interest are complex. This is particularly severe for the evaluation of the amplitudes for $B$-decays, where a large number of on-shell intermediate states can propagate and contribute to the imaginary part. In an attempt to overcome this complication, methods based on the spectral representation have been developed [49, 50, 51, 52, 23, 53, 54, 24, 55, 56] and applied to a variety of physical processes, including Refs. [57, 58, 59, 60, 61, 62, 63, 64, 47, 65, 66, 67, 68, 69]. In particular, in Ref. [24] a method has been proposed which allows for the determination of both the real and imaginary parts of complex amplitudes in which the hadronic factor is of the generic form:
@@ III.2 Time-ordered diagrams contributing to $\bar{B}_{s}\to\gamma\ell^{+}\ell^{-}$ decay amplitudes
Figure 2: Schematic illustrations of the two time-orderings in Eq. (11) contributing to $B\to K\ell^{+}\ell^{-}$ decays. The dashed lines between the operators represent all propagating states with the indicated $B$ and $S$ quantum numbers. The emission of the virtual photon is not shown. Panel (a) corresponds to the time ordering $t<0$, while panel (b) corresponds to $t>0$.
@@
Figure 3: Schematic illustrations of the three time-orderings in Eq. (14) which contribute an imaginary part to the $B\to\gamma\ell^{+}\ell^{-}$ decay amplitude. The dashed lines between the operators represent all propagating states with the indicated $B$ quantum number. The emission of the real and virtual photons is not shown. Panel (a) corresponds to the time ordering $t_{W}<0<t$, panel (b) to $t_{W}<t<0$, while panel (c) to $t<t_{W}<0$.
@@ III.3 Generalized spectral density representation
where the subscript $E$ on the left-hand side denotes Euclidean  [^6] Note that $H^{\nu+}_{1,2}(\vec{q}\hskip 1.2pt)=\int_{0}^{\infty}dt\,e^{q_{0}t}\,C_{1,2;E}^{\nu+}(t,\vec{q}\hskip 1.2pt)$ is an incorrect continuation of Eq. (29) to Euclidean space. This time integral diverges at large $t$ due to the contributions from intermediate states with energies smaller than $m_{B}$, thus preventing the physical amplitude from being obtained.. The HLT method [23] is based on the expansion of the smearing kernel in terms of exponentials as follows:
@@
The issues concerning the limit $\epsilon\to 0$ have been been discussed in detail in Ref. [24] and are sketched in Sec. VIII.2 below, when we discuss our numerical results. The first issue concerns the fact that the computations are performed on finite spatial volumes with discrete spectra requiring the condition $\epsilon L\gg 1$ to obtain reliable smeared results. When this condition is satisfied, finite volume effects, which are known to be exponentially small in $\epsilon L$ [53], can be safely kept under control. The second issue is that the smallest values of $\epsilon$ that can be reached in practice, depends on both the finite numbers of discrete lattice points at which the relevant time-dependent correlation function has been computed, i.e. on the temporal extension T of the lattice, and on the statistical accuracy of the lattice data. In order to ensure good control of the extrapolation $\epsilon\to 0$, at any selected value of $E$, the condition $\epsilon\ll\Delta(E)$ must be satisfied, where $\Delta(E)$ indicates the typical size of the interval around $E$ in which the hadronic amplitude is significantly varying, such as in a region where narrow resonances contribute to the spectral density. In such cases the condition $\epsilon\ll\Delta(E)$ may be difficult to be fulfilled. In such cases, in order to obtain results with sufficiently small values of $\epsilon$ will require extremely high statistical accuracy of the lattice data. For this reason, in our numerical study in Sec. VIII.2 we evaluate the amplitude at a value of $q^{2}$ which is above the squared masses of the narrow charmonium resonances, where the amplitude behaves relatively smoothly. In such an asymptotic regime, one can expect that the extrapolation can be performed using a polynomial ansatz.
@@ VII.1 The contact terms
The term $H^{\nu+;\,3\mathrm{subs}}_{1,2}(\vec{q},\epsilon)$, defined in Eq. (77), has been designed so that it has no ultraviolet divergences, even at finite-values of $\epsilon$. We envisage applying the HLT procedure to each of the four terms in Eq. (77). The first three terms on the right-hand side of Eq. (76) compensate for the subtractions introduced in the definition of $H^{\nu+;\,3\mathrm{subs}}_{1,2}(\vec{q},\epsilon)$ given in Eq. (77). They have no poles in the range of the integration over $E$ so that we can set $\epsilon=0$ in these terms. They do however, have quadratic and linear ultraviolet divergences. In order to cancel the power divergences, we combine these three terms with the contribution from the region $t<0$ in Eq. (IV) in a correlated way so that only the logarithmic divergence remains. In a lattice QCD computation this combination is performed through a jackknife analysis, with the expectation that the correlations between the four terms will reduce the resulting uncertainty.
@@
Figure 8: Panels (a) and (b) show the two diagrams for the decay $\bar{B}_{s}\to\mu^{+}\mu^{-}\gamma$ in which both the real photon $\gamma$ with momentum $k$ and the virtual photon with momentum $q$ are emitted from the charm-quark loop.
A corresponding procedure needs to be developed and applied to the contact terms present in the amplitude for $\bar{B}_{s}\to\gamma\ell^{+}\ell^{-}$ decays. We postpone the presentation of the details to a future publication, when we will be in a position to have learned from the experience of computing the charming penguin contributions to $B\to K\ell^{+}\ell^{-}$ decay amplitudes. Here we only note that for the process $\bar{B}_{s}\to\gamma\ell^{+}\ell^{-}$, there are also diagrams in which both the real and virtual photons are emitted from the charm-quark loop as in Fig. 8. Although the superficial degree of divergence of such diagrams would allow for a logarithmic ultraviolet divergence, and indeed, even at lowest order, each of the two diagrams in Fig. 8 separately has such a divergence, electromagnetic gauge invariance ensures that sum of the two contributions from these diagrams is finite. The requirement that the matrix element in Eq. (50) vanishes when contracted with $k^{\mu}$ and $q^{\nu}=(p_{\bar{B}_{s}}-k)^{\nu}$, implies that the matrix element is proportional to two powers of the external momenta which is sufficient to make it finite. However, similarly to the above discussion of the cancellation of the power divergences in $B\to K\ell^{+}\ell^{-}$ decays, the cancellation of the logarithmic ultraviolet divergences in the diagrams of Fig.8 requires the sum of the contributions from all the time-orderings. This is true since we use the lattice conserved vector current. Had we used the local vector current a more complicated subtraction procedure would need to be implemented but we do not discuss this here.
@@ VII.2.4 Renormalization of the operator $O_{8}$
Figure 9: Top Panel: Bare correlators $\widetilde{C}^{3}_{1,2;VV,AA}(t,\vec{q}\hskip 1.2pt)$, defined in Eqs. (129)-(130), corresponding to the $B\to K\ell^{+}\ell^{-}$ decay. Bottom: The correlators $\widetilde{C}^{3}_{1,2;VV,AA}(t,\vec{q}\hskip 1.2pt)$ corresponding to the $B\to\eta_{ss^{\prime}}\ell^{+}\ell^{-}$ decay, obtained by replacing the (spectator) light-quark in the $B\to K\ell^{+}\ell^{-}$ correlation function by a strange quark as described in the text. $N=1$ and $N=2$ correspond to the operators $O_{1}^{(c)}$ and $O_{2}^{(c)}$ respectively. In each case the results have been multiplied by $t^{12}$, i.e. $t$ to the power 12, for visualization purposes.
@@
$\widetilde{C}_{1,2;VV,AA}^{\nu}(t,\vec{q}\hskip 1.2pt)=\int_{E^{*}}^{\infty}\frac{dE}{2\pi}e^{-Et}\rho^{\nu+}_{1,2;VV,AA}(E,\vec{q}\hskip 1.2pt)~,\qquad\rho^{\nu+}_{1,2}(t,\vec{q}\hskip 1.2pt)=\rho^{\nu+}_{1,2;VV}(t,\vec{q}\hskip 1.2pt)+\rho^{\nu+}_{1,2;AA}(t,\vec{q}\hskip 1.2pt)~.$ (131)
@@ VIII.1 From the Euclidean correlators to the amplitude using the SFR/HLT method
where $\rm Cov$} is the covariance matrix of the Euclidean lattice correlator $\widetilde{C}_{1,2}^{\nu+}(t,\vec{q}\hskip 1.2pt;m_{H})$, and $\lambda$ is the so-called trade-off parameter. In Eq. (139), $\beta$ and $E_{\mathrm{th}}$ are algorithmic parameters which we set to $\beta=1.99$ and $E_{\rm th}=3.5\,{\rm GeV}$ ($3.7~{\rm GeV}$) in the case of $B\to K\ell^{+}\ell^{-}$ ($B_{s}\to\eta_{ss^{\prime}}\ell^{+}\ell^{-}$). [^14] For the convergence of the integral $\beta$ must be smaller than 2. In the presence of statistical errors, the functional $B$ disfavours coefficients $\bm{g}$ leading to large statistical uncertainties in the reconstructed smeared amplitude. The balance between having small systematic errors (small $A_{R}^{\beta}[\bm{g}]$ and $A_{I}^{\beta}[\bm{g}]$) and small statistical errors (small $B[\bm{g}]$) depends on the tunable parameter $\lambda$. Its optimal value $\lambda^{\rm opt}$ is determined by monitoring the evolution of the reconstructed smeared amplitude for different values of $\lambda$. The optimal value is chosen to be in the statistically-dominated regime, where $\lambda$ is sufficiently small that the systematic error due to the kernel reconstruction is smaller than the statistical error (therefore, in this region, the results are stable under variations of $\lambda$ within statistical uncertainties), but still large enough to have reasonably small statistical uncertainties. The kernel functions that we use at non-zero lattice spacing are obtained from Eq. (VIII.1) by the replacement